A total of 550 people attended a concert. Floor tickets cost $40 each, while balcony tickets cost $28 each. If a total of $20,800 was collected, how many of each type of ticket were sold?
Answer by checkley71(8403)

F+B=550 F=550-B
40F+28B=20,800 NOW REPLACE F WITH(550-B)
40(550-B)+28B=20,800
22,000-40B+28B=20,800
-12B=20,800-22,000
-12B=-1200
B=-1200/-12
B=100 TICKETS SOLD FOR THE BALCONY.
F=550-100
F=450 TICKETS SOLD FOR THE FLOOR.
PROOF
40*450+28*100=20,800
18,000+2,800=20,800
20,800=20,800